Solve the system by the method of elimination and check any solutions algebraically
{2x + 5y = 12
{5x + 8y = 15

(x, y) =

Answer by TimothyLamb(4379)

---
2x + 5y = 12
5x + 8y = 15
---
copy and paste the above standard form linear equations in to this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= -2.333333 = -7/3
y= 3.333333 = 10/3
---
---
now by elimination:
10x + 25y = 60
-10x - 16y = -30
---
add:
9y = 30
y = 10/3 = 3.333333
2x + 5(10/3) = 12
2x = 12 - 5(10/3)
2x = 36/3 - 50/3
2x = -14/3
x = -14/(2*3)
x = -7/3 = -2.333333
---
answer by elimination (same answer as by the matrix-method):
x= -2.333333 = -7/3
y= 3.333333 = 10/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php